OpenStudy (anonymous):

I need to integrate the function (6x+7y)^2 over the area of a triangle with vertices (-5,0), (0,5), (5,0). I'll post what I tried already; it was wrong.

OpenStudy (anonymous):

\[\int\limits_{-5}^{5} \int\limits_{0}^{5} (6x+7y)^2 dy dx\]This was wrong. I see now that this made a rectangular area.

OpenStudy (sirm3d):

\[\large \int\limits_{?}^{?}\int\limits_{R}^{?}(6x+7y)^2dA=\int\limits_{?}^{?}\int\limits_{R_1}^{?}(6x+7y)^2dA+\int\limits_{?}^{?}\int\limits_{R_2}^{?}(6x+7y)^2dA\]

OpenStudy (sirm3d):

\[\large =\int\limits_{-5}^{0}\int\limits_{0}^{(x+5)}(6x+7y)^2dy dx+\int\limits_{0}^{5}\int\limits_{0}^{(5-x)}(6x+7y)^2dy dx\]

OpenStudy (anonymous):

Thanks very much for that! I'll try it and see what I get. Very helpful for the 4 other questions I have to do.

OpenStudy (sirm3d):

YW

OpenStudy (sirm3d):

Here's a second solution: 

OpenStudy (sirm3d):

\[\large \int\limits_{}^{}\int\limits_{R}^{}(6x+7y)^2dA=\int\limits_{}^{}\int\limits_{R}^{}(6x+7y)^2dxdy\]\[\large =\int\limits_{y=0}^{y=5}\int\limits_{x=y-5}^{x=5-y}(6x+7y)^2dxdy\]